February 12, 2014 Calibration

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Wednesday January 15, 2013 at 11:00 AM. At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET

  1. ) ESNET: 8542553
  2. )To connect by telephone, dial:
  • You can look up toll-free number at http://www.readytalk.com/intl
  • US and Canada: (866)740-1260 (toll free)
  • International: (303)248-0285 (toll call)
  • enter access code followed by the # sign: 3421244#


  1. Announcements
  2. Status Update
  3. Tracking
  4. Calorimetry
  5. Other systems
    • Photon Beam
    • TOF
    • Start Counter
  6. Simulations and Calibrations


Attending: Sean, Amiran (NU); Simon, Elton, Mark, Beni, Lubomir, Eugene (JLab); Curtis, Paul (CMU); Zisis, Andrei (Regina); Justin (MIT).

  1. Announcements
    • Sean will give a talk at the CM summarizing the current status of calibrations.
  2. Status Update
    • Constants have been added to the CCDB for all subdetectors except for the Start Counter and Photon Beam. The next step is to identify manpower to update the reconstruction and simulation codes to use these.
  3. Tracking
    • Simon reported slow progress with tracking alignment.
    • Sean got a sample CDC time-to-distance relation from Naomi and added it to the CCDB for testing purposes. The plan is to make these relations for gas mixes between 40/60 and 60/40 and select the appropriate relation at run time.
  4. Calorimetry
    • BCAL - There was much discussion about storing values in different databases. The Regina group has put together a comprehensive list of calibrations and conditions variables. This list needs to be rationalized with the systems that are and will be used. Yi mentioned that there currently exists the translation table DB and JInventory DBs, which are currently being extended for use in detector tests in an ad-hoc way. The online working group is in charge of defining and implementing the databases that will be used for actual data taking, and it is not clear what the current status of this effort is. There may need to be dedicated meetings between the people in charge of the online DBs and the detector groups to work out some of these details. It was hoped that many of the details can be worked out at informal conversations at the Collab. Meeting.
    • FCAL - Elton reported that they have started to look at pedestals, both with the bases off and on.
  5. Other systems
    • Photon Beam - Richard is scheduled to report on this at the next meeting (2/26)
    • TOF - Work is ongoing based on test data and simulations using a standalone program.
    • Start Counter - Work is ongoing
  6. Simulations and Calibrations
    • Sean presented an initial look at rates from bggen for various channels which can be used as the input for calibration procedures.
    • The studies were done assuming a commissioning beam of 10^6 photons/s, but assuming the photon spectrum from a diamond radiator. A smoother spectrum should be used.
    • It's not clear what the commissioning trigger will be like, but it was decided to only consider particles with at least 100 MeV in momentum/energy.
    • Elton had done an earlier look at rates of pi0 and eta decays into two photons (linked above). eta's should be considered and these results should be checked against his numbers.
    • Curtis asked Zisis if he was interested in events containing multiple pi0's. Zisis said his preference was to start with cleaner events with single pi0s, and then consider adding more once we know the detector better.
    • Justin suggested looking at states that decay into kaons to test PID performance. phi -> K+ K- will be looked at, K* -> K pi is also a possibility (if a bit wide).
    • Sean will refine this analysis by analyzing REST files, taking the above into account, as preparations for the data challenge progress.